Enter the value for x that makes the equation 23(x−6)=34(x+16) true.
Answer by richwmiller(17219)

23(x−6)=34(x+16)
23*x−6*23=34*x+16*34
subtract 23*x from both sides
multiply out 6*23 and 16*34
subtract product of 16*34 from both sides